# Approximating Quantum States by States of Low Rank

### v1

@@ Abstract
Given a positive integer $k$, it is natural to ask for a formula for the distance between a given density matrix (i.e., mixed quantum state) and the set of density matrices of rank at most $k$. This problem has already been solved when “distance” is measured in the trace or Frobenius norm. We solve it for all other unitary similarity invariant norms. We also present some consequences of our formula. For example, in the trace and Frobenius norms, the density matrix that is farthest from the set of low-rank density matrices is the maximally mixed state, but this is not true in many other unitary similarity invariant norms.
@@ 1 Introduction
We then explore several consequences of our result, such as the question of which quantum state is farthest from the set of low-rank quantum states. We show that the answer to this question is the maximally-mixed state in several cases, such as when the norm being used is the trace norm (Proposition 3.1). However, for other unitary similarity invariant norms, all that can be said in general is that the farthest state is one of nkn-k different candidates (Theorem 3.1), one of which is the maximally-mixed state.
We then specialize to the Schatten and Ky Fan norms. We show that if $p\in\{1\}\cup[2,4]$, then the farthest state when using the Schattne $p$-norm is always the maximally mixed state, but for all other values of $p$ there exists a rank $k$ and a dimension $n$ for which the farthest (in Schatten $p$-norm) state is not the maximally-mixed state (Theorem 3.2). We also determine the farthest state in the Ky Fan norms in Theorem 3.3.
@@ 3 The farthest state from the set of low-rank states
Now that we know how to find the closest low-rank state to a given state, we explore the question of which state is farthest from the set of low-rank states. We might expect $d(X,{\mathcal{D}}_{n,k})$ to be maximized when $X=\frac{1}{n}I_{n}\in{\mathcal{D}}_{n}$ (i.e., when $X$ is the maximally mixed state). It turns out that, for many unitary similarity invariant norms, this is not the case. For example, Theorem 2.1 tells us that if $n=4$, $k=3$, $\|\cdot\|$ is the operator norm, and $X=\mathrm{diag}\big(\frac{1}{3},\frac{1}{3},\frac{1}{3},0\big)$, then $d(X,{\mathcal{D}}_{n,k})=\frac{1}{3}>\frac{1}{4}=d(\tfrac{1}{n}I_{n},{\mathcal{D}}_{n,k})$.
@@
The situation for general USI norms is somewhat more complicated than it is for the trace norm. Our main result of this section shows that, even though $d(X,{\mathcal{D}}_{n,k})$ is not necessarily maximized by $X=\tfrac{1}{n}I_{n}$, it is always maximized by some state of the form $X=\frac{1}{m}I_{m}\oplus O_{n-m}$, where $m\in\{k+1,m+2,\ldots,n\}$:
@@
Now, if $x_{k}>x_{k+1}$ then we may change $(x_{k},x_{k+1})$ to $(x_{k}-d,x_{k+1}+d)$ for $d=\frac{1}{2}(x_{k}-x_{k+1})>0$ to get a matrix $\hat{X}$ (with best rank-$k$ approximation $\hat{Y}\in{\mathcal{D}}_{n}$) for which $\|X-Y\|\leq\|\hat{X}-\hat{Y}\|$. This follows from the fact that, if we define $\hat{\gamma}$ analogously to how we defined $\gamma$, but coming from the diagonal entries of $\hat{X}$ instead of $X$, then the vector $(-\gamma,\ldots,-\gamma,x_{k+1},x_{k+2},\ldots,x_{n})$ is majorized by $(-\hat{\gamma},\ldots,-\hat{\gamma},x_{k+1}+d,x_{k+2},\ldots,x_{n})$. It follows that that the maximum is attained when $X$ has the form $X={\rm diag}(x,\ldots,x,x,x_{k+2},\ldots,x_{n})$ for some real $x$.
@@
Finally, define $X_{1}=\frac{1}{m}I_{m}\oplus O_{n-m}$ and $X_{2}=\frac{1}{m+1}I_{m+1}\oplus O_{n-(m+1)}$. Then $Y=\frac{1}{k}I_{k}\oplus O_{n-k}\in{\mathcal{D}}_{n}$ is the best rank-$k$ approximation for each of $X$, $X_{1}$, and $X_{2}$. Since, XYX-Y is a convex combination of $X_{1}-B$ and $X_{2}-B$, we have
@@ 3.2 The Ky Fan norms
We now explore the question of which state $X\in{\mathcal{D}}_{n}$ is farthest from the set of rank-at-most-$k$ states when we measure distance via one of the Ky Fan norms. Theorem 3.1 of course applies in this setting and tells us that the farthest state is one of nkn-k different candidates, corresponding to the nk-n-k different choices of $m\in\{k+1,k+2,\ldots,n\}$. It follows that we could always just check which of those nkn-k different candidates is farthest from ${\mathcal{D}}_{n,k}$.

### v2

@@ Abstract
Given a positive integer $k$, it is natural to ask for a formula for the distance between a given density matrix (i.e., mixed quantum state) and the set of density matrices of rank at most $k$. This problem has already been solved when “distance” is measured in the trace or Frobenius norm. We solve it for all other unitary similarity invariant norms. We also present some consequences of our formula. For example, in the trace and Frobenius norms, the density matrix that is farthest from the set of low-rank density matrices is the maximally-mixed state, but this is not true in many other unitary similarity invariant norms.
@@ 1 Introduction
There are numerous potential applications of this main result. For example, low-rank quantum states can be stored and transmitted more efficiently than high-rank quantum states, since a rank-$k$ state can be specified by just $O(kn)$ real parameters (via its spectral decomposition) rather than the $n^{2}-1$ real parameters that are needed to specify an arbitrary state. It thus might be desirable in some limited-resource situations to approximate a quantum state by the nearest low-rank state and work with or transmit that state instead. As another example, in quantum information theory the problem of reconstructing a quantum state based on measurement results is called state tomography. Performing state tomography to reconstruct pure (i.e., rank-$1$) states is a task that has been thoroughly studied [CDJ+13, VD23], as has this task for other low ranks [AKG16]. Because measurements are noisy in practice, tomography often results in a quantum state that is not actually pure even though it should be, and the same is true when tomography is applied to states of other low ranks. Our results provide a way of converting the (incorrect) tomographically-reconstructed state into the nearest state that has the desired rank.
We then explore several consequences of our result, such as the question of which quantum state is farthest from the set of low-rank quantum states, or equivalently what the maximum possible distance between a quantum state and the set of low-rank quantum states is. Distance measures like this are common in the theory of quantum resources [REG18], and it is important to know what values they can take on.
For example, the well-studied geometric measure of entanglement asks for the distance between a bipartite pure state and the set of separable pure states, and this quantity is maximized exactly by the maximally-entangled states [WG03] (i.e., the states with all of their Schmidt coefficients equal to each other). We show that, similarly, the answer of the analogous question of which state is farthest from low-rank is the maximally-mixed state in several cases, such as when the norm being used is the trace norm (Proposition 3.1). However, we also show that there are many other unitary similarity invariant norms for which this is not true. In general, all that can be said is that the farthest state is one of nkn-k different candidates (Theorem 3.1), one of which is the maximally-mixed state.
We then further investigate these questions for the Schatten and Ky Fan norms, since they are the most commonly-occurring families of unitary similarity-invariant norms in quantum information theory (see [WAT05, WAT18, HO22] and the references therein, for example). The Schatten norms occur frequently in this area because the trace, Frobenius, and operator norms occur as special cases of them, and because of their connection to the (now-disproved) additivity conjecture [HW08, HAS09]. The Ky Fan norms similarly make frequent appearances in the field because of their connection to majorization, which is absolutely ubiquitous in the field [NIE02, JAB13].
In particular, we show that if $p\in\{1\}\cup[2,4]$, then the farthest state from the set of low-rank states when using the Schatten $p$-norm is always the maximally-mixed state, but for all other values of $p$ there exists a rank $k$ and a dimension $n$ for which the farthest state (in Schatten $p$-norm) is not the maximally-mixed state (Theorem 3.2). We also determine the farthest state in the Ky Fan norms in Theorem 3.3.
@@ 3 The farthest state from the set of low-rank states
Now that we know how to find the closest low-rank state to a given state, we explore the question of which state is farthest from the set of low-rank states. We might expect $d(X,{\mathcal{D}}_{n,k})$ to be maximized when $X=\frac{1}{n}I_{n}\in{\mathcal{D}}_{n}$ (i.e., when $X$ is the maximally-mixed state). It turns out that, for many unitary similarity invariant norms, this is not the case. For example, Theorem 2.1 tells us that if $n=4$, $k=3$, $\|\cdot\|$ is the operator norm, and $X=\mathrm{diag}\big(\frac{1}{3},\frac{1}{3},\frac{1}{3},0\big)$, then $d(X,{\mathcal{D}}_{n,k})=\frac{1}{3}>\frac{1}{4}=d(\tfrac{1}{n}I_{n},{\mathcal{D}}_{n,k})$.
@@
The situation for general USI norms is somewhat more complicated than it is for the trace norm. Our main result of this section shows that, even though $d(X,{\mathcal{D}}_{n,k})$ is not necessarily maximized by $X=\tfrac{1}{n}I_{n}$, it is always maximized by some state of the form $X=\frac{1}{m}I_{m}\oplus O_{n-m}$, where $m\in\{k+1,k+2,\ldots,n\}$:
@@
Now, if $x_{k}>x_{k+1}$ then we may change $(x_{k},x_{k+1})$ to $(x_{k}-d,x_{k+1}+d)$ for $d=\frac{1}{2}(x_{k}-x_{k+1})>0$ to get a matrix $\hat{X}$ (with best rank-$k$ approximation $\hat{Y}\in{\mathcal{D}}_{n}$) for which $\|X-Y\|\leq\|\hat{X}-\hat{Y}\|$. This follows from the fact that, if we define $\hat{\gamma}$ analogously to how we defined $\gamma$, but coming from the diagonal entries of $\hat{X}$ instead of $X$, then the vector $(-\gamma,\ldots,-\gamma,x_{k+1},x_{k+2},\ldots,x_{n})$ is majorized by $(-\hat{\gamma},\ldots,-\hat{\gamma},x_{k+1}+d,x_{k+2},\ldots,x_{n})$. It follows that the maximum is attained when $X$ has the form $X={\rm diag}(x,\ldots,x,x,x_{k+2},\ldots,x_{n})$ for some real $x$.
@@
Finally, define $X_{1}=\frac{1}{m}I_{m}\oplus O_{n-m}$ and $X_{2}=\frac{1}{m+1}I_{m+1}\oplus O_{n-(m+1)}$. Then $Y=\frac{1}{k}I_{k}\oplus O_{n-k}\in{\mathcal{D}}_{n}$ is the best rank-$k$ approximation for each of $X$, $X_{1}$, and $X_{2}$. Since XYX-Y is a convex combination of $X_{1}-Y$ and $X_{2}-Y$, we have
@@ 3.2 The Ky Fan norms
We now explore the question of which state $X\in{\mathcal{D}}_{n}$ is farthest from the set of rank-at-most-$k$ states when we measure distance via one of the Ky Fan norms. Theorem 3.1 of course applies in this setting and tells us that the farthest state is one of nkn-k different candidates, corresponding to the nkn-k different choices of $m\in\{k+1,k+2,\ldots,n\}$. It follows that we could always just check which of those nkn-k different candidates is farthest from ${\mathcal{D}}_{n,k}$.